# ImagineNav++: Prompting Vision-Language Models as Embodied Navigator through Scene Imagination

### v1

@@ Abstract
Visual navigation is a fundamental capability for autonomous home-assistance robots, enabling the execution of long-horizon tasks such as object search. While recent methods have leveraged Large Language Models (LLMs) to incorporate commonsense reasoning and improve exploration efficiency, their planning processes remain constrained by textual representations, which cannot adequately capture spatial occupancy or scene geometry–critical factors for informed navigation decisions. In this work, we explore whether Vision-Language Models (VLMs) can achieve mapless visual navigation using only onboard RGB/RGB-D streams, unlocking their potential for spatial perception and planning. We achieve this by developing the imagination-powered navigation framework ImagineNav++, which imagines the future observation images at valuable robot views and translates the complex navigation planning process into a rather simple best-view image selection problem for VLMs. Specifically, we first introduce a future-view imagination module, which distills human navigation preferences to generate semantically meaningful candidate viewpoints with high exploration potential. These imagined future views then serve as visual prompts for the VLM to identify the most informative viewpoint. To maintain spatial consistency, we develop a selective foveation memory mechanism, which hierarchically integrate keyframe observations through a sparse-to-dense framework, thereby constructing a compact yet comprehensive memory for long-term spatial reasoning. This integrated approach effectively transforms the challenging goal-oriented navigation problem into a series of tractable point-goal navigation tasks. Extensive experiments on open-vocabulary object and instance navigation benchmarks demonstrate that our ImagineNav++ achieves SOTA performance in mapless setting, even surpassing most cumbersome map-based methods, revealing the importance of scene imagination and scene memory in VLM-based spatial reasoning.
@@ I Introduction
A useful home-assistant robot should be able to search for different kinds of objects without access to precise 3D spatial coordinates. This capability becomes particularly essential given the dynamic nature of household environments, where new objects are continuously introduced through regular human acquisition patterns. Consequently, the robot’s navigation system requires open-set visual grounding and searching functionality that extends beyond predefined categorical constraints. In the research community, this fundamental challenge is formally characterized as the open-vocabulary goal-oriented visual navigation problem. Depending on the type of goal specification, several sub-tasks have been studied within this paradigm. Among these, we focus on two representative categories: Object-Goal Navigation (ObjectNav), in which the target is specified by a semantic category [topiwala2018frontier, cai2023bridgingzeroshotobjectnavigation, chaplot2020object, ramrakhya2022habitat], and Instance-Image-Goal Navigation (InsINav), where the goal is defined by a reference image of a specific object instance [sun2024prioritizedsemanticlearningzeroshot, yin2025unigoaluniversalzeroshotgoaloriented].
@@
Recent advances in foundation models, including vision models [radford2021learning, he2022masked, zhou2022detecting, Cheng2024YOLOWorld, kirillov2023segment, wu2024general, liu2023grounding], large language models (LLMs) [brown2020language, chowdhery2023palm, touvron2023llama, zhang2022opt], and vision-language models (VLMs) [achiam2023gpt, team2023gemini, liu2024visual, chen2024internvl, Dai2023InstructBLIPTG, gao2023llama], have enabled new possibilities for developing agents capable of open-vocabulary goal-oriented visual navigation. A widely adopted framework, as illustrated in Figure 1, employs a modular approach to address this challenge, typically comprising four key components: (1) A real-time mapping and segmentation module to constructs a geometric and semantic representation of the robot’s immediate environment through simultaneous localization and mapping (SLAM) combined with pixel-wise semantic segmentation [he2018maskrcnn, kirillov2023segment]; (2) A template-based translation module to transform the structured semantic map into natural language descriptions, enabling compatibility with language models; (3) A LLM-based reasoning module to generate step-by-step navigation plans in texts, incorporating task constraints and commonsense reasoning; (4) Finally, a path-planning module to projects the LLM’s output back onto the environmental map, computing optimal collision-free trajectories towards the goal using off-the-shelf motion planning algorithms. This pipeline effectively combines the strengths of visual perception, language understanding, and robotic control to achieve open-vocabulary navigation capabilities.
Although such pipelines have achieved significant success in recent years [Zhou2023ESCEW, kuang2024openfmnav, wu2024voronav, zhang2024trihelper, shah2023navigationlargelanguagemodels, Yu2023L3MVNLL, loo2025open], these cascaded systems face several inherent limitations. First, both the depth camera and the robot localization module can suffer from perception error, especially for long-range depth estimation, and this can make the mapping process inaccurate. Second, the robotic system must maintain continuous real-time object detection and semantic segmentation to simultaneously enrich spatial mapping with semantic metadata and generate structured inputs for LLM-based reasoning. These computationally intensive operations introduce a critical performance bottleneck, substantially elevating the processing requirements of the robotic platforms. Third, while semantic information stored in the map can be readily represented in textual form–for instance, by listing observed object categories–such text-only prompts struggle to convey explicit geometric relationships and fine-grained object details. This limitation makes it difficult and ambiguous for LLMs to infer the best navigation plan.
@@
Furthermore, it is well-established that agent memory plays a crucial role in embodied navigation, allowing agents to accumulate, interpret, and leverage historical scene observations to inform future decisions [fukushima2022object, wang2023gridmm]. A major open challenge that arises in the context of VLM-based view selection is how to effectively encode the continuously expanding stream of observations to endow pre-trained VLMs with temporal reasoning capabilities. A common approach leverages VLMs to caption individual observations, subsequently aggregating these textual summaries over time [zhou2023navgpt, zhang2025mem2ego]. However, this abstraction process may inadvertently discard essential spatial and semantic information inherent in raw visual data, fundamentally limiting effective spatio-temporal reasoning. An alternative paradigm directly employs historical image sequences for end-to-end decision-making [zhang2024navid, zhang2024uni, tsai2023multimodal]. While architecturally simpler, these approaches often exhibit limited capacity to model long-term dependencies and show reduced robustness against perceptual noise in lengthy visual sequences. To overcome these limitations, we develop a Selective Foveation Memory, which mimics human foveation mechanisms by hierarchically integrating keyframe observations through a sparse-to-dense framework–maintaining sparse keyframes for long-range structural context while preserving dense frames for proximate details, thereby efficiently capturing both global layout and local semantics. In particular, we introduce the innovative use of the pre-trained vision foundation model DINOv2 [DINOv2] to semantically interpret historical observations, measure inter-frame similarity, and extract representative keyframes from extended observation sequences, effectively addressing the redundancy in long-term visual streams while preserving essential spatio-temporal structures for robust decision-making.
@@
  • For zero-shot ObjectNav, our ImagineNav++ increases success rate by a large margin of 4.0% and 23.5% respectively on the complex HM3D [ramakrishnan2021hm3dc] and HSSD [khanna2023habitatsyntheticscenesdataset], while achieving competitive result on Gibson [xia2018gibson]. For InsINav, our approach attains the highest SPL on HM3D.
This paper is an extended version of our prior publication [zhao2025imaginenavpromptingvisionlanguagemodels] in ICLR 2025. The main differences from the conference version are listed as follows: 1) We introduce a selective foveation memory that maintains spatial consistency via structured trajectory reasoning, thereby mitigating prior decision biases through its hierarchical keyframe selection and organization. 2) We demonstrate the method’s robust transferability by showing that it could seamlessly adapt to the InsINav task with only minimal architectural adjustments. 3) Our empirical findings are further supported by extensive evaluation across multiple benchmarks, complemented by detailed ablation studies and hyperparameter analysis.
@@ II-A Large Models for Robotic Planning
Large-scale models pre-trained on extensive internet data have demonstrated formidable zero-shot reasoning capabilities in tasks such as planning [huang2022languagemodelszeroshotplanners], code generation [liang2023codepolicieslanguagemodel, huang2023instruct2act], and solving science questions [lewkowycz2022solvingquantitativereasoningproblems]. The in-context learning capability of LLMs allows them to be applied to robotic task planning. Some methods [liang2023codepolicieslanguagemodel, huang2023instruct2act, ahn2022icanisay] leverage LLMs to decompose tasks into subtasks, enhancing execution efficiency. Cap [liang2023codepolicieslanguagemodel] generates robotic policy code directly from example language commands, enabling autonomous control and task execution based on natural language instructions. Instruct2Act [huang2023instruct2act] combines LLM with foundational models (e.g., SAM and CLIP), reducing error rates in complex task execution, while SayCan [ahn2022icanisay] combines LLM task planning with the feasibility of physical skills using pre-trained value functions, generating actionable plans for robots. However, one limitation of LLMs is their difficulty in embedding the robot’s state directly into the planning process. To address this, many studies have turned to VLMs as alternatives. For instance, ViLA [lin2024vilapretrainingvisuallanguage] significantly improves performance on multimodal tasks without compromising text capabilities by systematically exploring VLM pretraining design choices. CoPa [huang2024copageneralroboticmanipulation] incorporates commonsense knowledge from VLMs, proposing a coarse-to-fine task-oriented grasping and task-aware motion planning approach. PIVOT [nasiriany2024pivotiterativevisualprompting] transforms tasks into iteratively optimized visual question-answering problems via a refinement process. Socratic model [zeng2022socraticmodelscomposingzeroshot] integrates multiple pretrained large models (e.g., VLMs, LLMs, and audio models) in a modular fashion to enable reasoning and task execution through language-based interaction. These methods employ a set-of-examples (SOE) approach to guide VLM selection. We propose a new decision-making paradigm based on imagined imagery, wherein decisions are made on imaginations, enabling more nuanced, context-aware interactions that better harness VLMs’ spatial perception capabilities.
@@ II-B Open-Vocabulary Navigation: From Objects to Instances
Open-vocabulary navigation requires agents to respond to instructions involving object categories or specific instances not encountered during training, placing higher demands on the model’s generalization and grounding capabilities. In object goal navigation, early end-to-end methods attempted to integrate textual knowledge into navigation tasks by leveraging compact multimodal feature spaces such as CLIP [radford2021learning], yet their performance remains limited [khandelwal2022simple, gadre2023cows, majumdar2022zson]. Modular approaches [huang2023visuallanguagemapsrobot, Zhou2023ESCEW, achiam2023gpt], on the other hand, typically rely on sensors for localization and mapping, high-level planning, and low-level control. Their dependence on hardware precision and pre-computation constrains practical flexibility. Meanwhile, instance goal navigation presents greater challenges due to its requirement for distinguishing unique target instances. Representative works such as Goat [chang2024goatthing] achieve lifelong navigation by constructing an instance-aware semantic memory and integrating classical path planning. Mod-IIN [krantz2023navigatingobjectsspecifiedimages], is specifically designed for this task, utilizing feature matching and map projection for goal re-localization. UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] employs a unified graph representation and graph matching to handle diverse goal types. While these methods demonstrate strong performance on specific benchmarks, most still rely heavily on explicit environmental representations–such as maps or scene graphs–or extensive task-specific training. Our approach introduces an imagination-based, mapless navigation framework. This framework circumvents the need for extensive training by transforming the complex process of navigation planning into a selection problem based solely on RGB inputs.
Figure 2: The overall pipeline of our mapless, open-vocabulary navigation framework ImagineNav++. At each iteration, the agent captures a panoramic view of its surroundings. The Imagination Module then leverage the trained Where2Imagine module coupled with a novel view synthesis model to generate novel scene views. Guided by structured prompts, the VLM engages in target-oriented inference by integrating historical selective foveation memory with imagined future waypoint observations. Subsequently, the system executes the PointNav policy to determine the next navigational action. The above imagination, reasoning and planning procedure iterates until the target is reached.
@@ II-C Imagination in Embodied Navigation
Recent methods [zhai2022peanutpredictingnavigatingunseen, ramakrishnan2022poni, zhu2022navigatingobjectsunseenenvironments] have adopted supervised learning to infer potential target locations, addressing the ‘Where to look?’ challenge in navigation. These approaches typically predict either the absolute coordinates [zhai2022peanutpredictingnavigatingunseen] of the target, the shortest distance to target [ramakrishnan2022poni], or the nearest boundary [zhu2022navigatingobjectsunseenenvironments] based on local maps. Another line of research [ramakrishnan2020occupancyanticipationefficientexploration, georgakis2021learning, 9560925, Zhang_2024_CVPR] has focused on enhancing the prediction of unobserved regions through diverse methodologies. For instance, [ramakrishnan2020occupancyanticipationefficientexploration] introduced occupancy anticipation, where the agent infers an occupancy map based on RGB-D inputs. The L2M framework was introduced [georgakis2021learning], consisting of a two-stage segmentation model that generates a semantic map beyond the agent’s field of view and selects long-term goals based on the uncertainty of predictions. SSCNav [9560925] leverages semantic scene completion and confidence maps to infer the environment and guide navigation decisions. A self-supervised generative map (SGM) is proposed [Zhang_2024_CVPR], which employs self-supervised approach to continually generate unobserved regions in the local map and predict the target’s location. These methods primarily predict unobserved regions in top-down maps derived from egocentric RGB-D projections. By contrast, our ImagineNav++ performs future-scene imagination directly in RGB space. We employ a compact model aligned with human navigation behavior to generate novel viewpoint locations, which are then translated into corresponding visual observations through a diffusion-based view synthesis process.
@@ II-D Memory Mechanism in Embodied Navigation
Memory plays a crucial role in visual navigation by improving an agent’s perceptual consistency and long-term reasoning capabilities in unfamiliar environments [fukushima2022object, wang2023gridmm]. Initially, LLMs were predominantly applied in embodied navigation systems to support high-level decision-making. To enable LLMs to process perceptual inputs, earlier approaches [zhou2023navgpt, zhang2025mem2ego] generate descriptive captions of individual observations and aggregate these textual summaries over time. Subsequent methods typically construct a globally consistent memory map from visual observations, infer potential navigational waypoints through structured spatial reasoning, and encode these into natural language representations suitable for LLMs [shah2023navigationlargelanguagemodels, wu2024voronav, kuang2024openfmnav, Zhou2023ESCEW], such as the semantic score map in OpenFMNav [kuang2024openfmnav] and the reduced Voronoi graph in VoroNav [wu2024voronav]. The inherent loss of geometric information in language representations has motivated the development of hybrid memory approaches that maintain structured environmental state values (e.g., target distance and exploration reward). For instance, VLMap [huang2023visuallanguagemapsrobot] quantifies navigation value via pixel-level feature-text similarity, while VLFM [yokoyama2024vlfm] projects image-text similarity into dynamic value maps using BLIP-2 [li2023blip]. While effectively mitigating memory forgetting, these methods require extensive precomputation and environment-specific training, limiting their scalability to unseen environments. With recent advances in multimodal learning, VLMs have been increasingly integrated into visual navigation systems to improve scene interpretation and decision-making [goetting2024end, tang2025openin, gao2025octonav, zhang2024navid, zhang2024uni, tsai2023multimodal]. VLMNav [goetting2024end] constructs dynamic traversability masks in a VQA framework. Concurrently, mapless navigation paradigms have emerged, bypassing explicit mapping to leverage multimodal observations. Representative methods include NaVid [zhang2024navid], which encodes monocular RGB streams into spatio-temporal representations using special tokens to differentiate historical and current observations; its extension Uni-NaVid [zhang2024uni] incorporating hierarchical token compression for efficient long-context processing; and the GRU-based history aggregation approach of [tsai2023multimodal], which models temporal observations to predict collision-aware action distributions. In line with these mapless navigation paradigms, our method also utilizes historical visual observations as direct input to VLMs. However, a key distinction of our approach lies in its use of visually salient keyframes, which compress perceptual history, enhance temporal coherence, and emphasize task-critical visual information, leading to more robust and efficient navigation performance.

### v2

@@ Abstract
Visual navigation is a fundamental capability for autonomous home-assistance robots, enabling the execution of long-horizon tasks such as object search. While recent methods have leveraged Large Language Models (LLMs) to incorporate commonsense reasoning and improve exploration efficiency, their planning processes remain constrained by textual representations, which cannot adequately capture spatial occupancy or scene geometry–critical factors for informed navigation decisions. In this work, we explore whether Vision-Language Models (VLMs) can achieve mapless visual navigation using only onboard RGB/RGB-D streams, unlocking their potential for spatial perception and planning. We achieve this by developing the imagination-powered navigation framework ImagineNav++, which imagines the future observation images at valuable robot views and translates the complex navigation planning process into a rather simple best-view image selection problem for VLMs. Specifically, we first introduce a future-view imagination module, which distills human navigation preferences to generate semantically meaningful candidate viewpoints with high exploration potential. These imagined future views then serve as visual prompts for the VLM to identify the most informative viewpoint. To maintain spatial consistency, we develop a selective foveation memory mechanism, which hierarchically integrates keyframe observations through a sparse-to-dense framework, thereby constructing a compact yet comprehensive memory for long-term spatial reasoning. This integrated approach effectively transforms the challenging goal-oriented navigation problem into a series of tractable point-goal navigation tasks. Extensive experiments on open-vocabulary object and instance navigation benchmarks demonstrate that our ImagineNav++ achieves SOTA performance in mapless setting, even surpassing most cumbersome map-based methods, revealing the importance of scene imagination and scene memory in VLM-based spatial reasoning.
@@ I Introduction
A useful home-assistant robot should be able to search for different kinds of objects without access to precise 3D spatial coordinates. This capability becomes particularly essential given the dynamic nature of household environments, where new objects are continuously introduced through regular human acquisition patterns. Consequently, the robot’s navigation system requires open-set visual grounding and searching functionality that extends beyond predefined categorical constraints. In the research community, this fundamental challenge is formally characterized as the open-vocabulary goal-oriented visual navigation problem. Depending on the type of goal specification, several sub-tasks have been studied within this paradigm. Among these, we focus on two representative categories: Object-Goal Navigation (ObjectNav), in which the target is specified by a semantic category [67, 6, 10, 57], and Instance-Image-Goal Navigation (InsINav), where the goal is defined by a reference image of a specific object instance [64, 84].
@@
Recent advances in foundation models, including vision models [53, 23, 101, 12, 33, 75, 45], large language models (LLMs) [5, 13, 68, 96], and vision-language models (VLMs) [1, 66, 44, 11, 14, 20], have enabled new possibilities for developing agents capable of open-vocabulary goal-oriented visual navigation. A widely adopted framework, as illustrated in Figure 1, employs a modular approach to address this challenge, typically comprising four key components: (1) A real-time mapping and segmentation module to construct a geometric and semantic representation of the robot’s immediate environment through simultaneous localization and mapping (SLAM) combined with pixel-wise semantic segmentation [24, 33]; (2) A template-based translation module to transform the structured semantic map into natural language descriptions, enabling compatibility with language models; (3) A LLM-based reasoning module to generate step-by-step navigation plans in texts, incorporating task constraints and commonsense reasoning; (4) Finally, a path-planning module to projects the LLM’s output back onto the environmental map, computing optimal collision-free trajectories towards the goal using off-the-shelf motion planning algorithms. This pipeline effectively combines the strengths of visual perception, language understanding, and robotic control to achieve open-vocabulary navigation capabilities.
Although such pipelines have achieved significant success in recent years [100, 36, 76, 94, 63, 86, 46], these cascaded systems face several inherent limitations. First, both the depth camera and the robot localization module can suffer from perception error, especially for long-range depth estimation, and this can make the mapping process inaccurate. Second, the robotic system must maintain continuous real-time object detection and semantic segmentation to simultaneously enrich spatial mapping with semantic metadata and generate structured inputs for LLM-based reasoning. These computationally intensive operations introduce a critical performance bottleneck, substantially elevating the processing requirements of the robotic platforms. Third, while semantic information stored in the map can be readily represented in textual form–for instance, by listing observed object categories–such text-only prompts struggle to convey explicit geometric relationships and fine-grained object details. This limitation makes it difficult and ambiguous for LLMs to infer the best navigation plan.
@@
Furthermore, it is well-established that agent memory plays a crucial role in embodied navigation, allowing agents to accumulate, interpret, and leverage historical scene observations to inform future decisions [17, 71]. A major open challenge that arises in the context of VLM-based view selection is how to effectively encode the continuously expanding stream of observations to endow pre-trained VLMs with temporal reasoning capabilities. A common approach leverages VLMs to caption individual observations, subsequently aggregating these textual summaries over time [99, 93]. However, this abstraction process may inadvertently discard essential spatial and semantic information inherent in raw visual data, fundamentally limiting effective spatio-temporal reasoning. An alternative paradigm directly employs historical image sequences for end-to-end decision-making [92, 91, 69]. While architecturally simpler, these approaches often exhibit limited capacity to model long-term dependencies and show reduced robustness against perceptual noise in lengthy visual sequences. To overcome these limitations, we develop a Selective Foveation Memory, which mimics human foveation mechanisms by hierarchically integrating keyframe observations through a sparse-to-dense framework–maintaining sparse keyframes for long-range structural context while preserving dense frames for proximate details, thereby efficiently capturing both global layout and local semantics. In particular, we introduce the innovative use of the pre-trained vision foundation model DINOv2 [51] to semantically interpret historical observations, measure inter-frame similarity, and extract representative keyframes from extended observation sequences, effectively addressing the redundancy in long-term visual streams while preserving essential spatio-temporal structures for robust decision-making.
@@
  • For zero-shot ObjectNav, our ImagineNav++ increases success rate by a large margin of 4.0% and 23.5% respectively on the complex HM3D [56] and HSSD [32], while achieving competitive result on Gibson [77]. For InsINav, our approach attains the highest SPL on HM3D.
This paper is an extended version of our prior publication [98] in ICLR 2025. The main differences from the conference version are listed as follows: 1) We introduce a selective foveation memory that maintains spatial consistency via structured trajectory reasoning, thereby mitigating prior decision biases through its hierarchical keyframe selection and organization. 2) We demonstrate the method’s robust transferability by showing that it could seamlessly adapt to the InsINav task with only minimal architectural adjustments. 3) Our empirical findings are further supported by extensive evaluation across multiple benchmarks, complemented by detailed ablation studies and hyperparameter analysis.
@@ II-A Large Models for Robotic Planning
Large-scale models pre-trained on extensive internet data have demonstrated formidable zero-shot reasoning capabilities in tasks such as planning [29], code generation [39, 28], and solving science questions [37]. The in-context learning capability of LLMs allows them to be applied to robotic task planning. Some methods [39, 28, 2] leverage LLMs to decompose tasks into subtasks, enhancing execution efficiency. Cap [39] generates robotic policy code directly from example language commands, enabling autonomous control and task execution based on natural language instructions. Instruct2Act [28] combines LLM with foundational models (e.g., SAM and CLIP), reducing error rates in complex task execution, while SayCan [2] combines LLM task planning with the feasibility of physical skills using pre-trained value functions, generating actionable plans for robots. However, one limitation of LLMs is their difficulty in embedding the robot’s state directly into the planning process. To address this, many studies have turned to VLMs as alternatives. For instance, ViLA [42] significantly improves performance on multimodal tasks without compromising text capabilities by systematically exploring VLM pretraining design choices. CoPa [27] incorporates commonsense knowledge from VLMs, proposing a coarse-to-fine task-oriented grasping and task-aware motion planning approach. PIVOT [48] transforms tasks into iteratively optimized visual question-answering problems via a refinement process. Socratic model [89] integrates multiple pretrained large models (e.g., VLMs, LLMs, and audio models) in a modular fashion to enable reasoning and task execution through language-based interaction. These methods employ a set-of-examples (SOE) approach to guide VLM selection. We propose a new decision-making paradigm based on imagined imagery, wherein decisions are made on imaginations, enabling more nuanced, context-aware interactions that better harness VLMs’ spatial perception capabilities.
@@ II-B Open-Vocabulary Navigation: From Objects to Instances
Open-vocabulary navigation requires agents to respond to instructions involving object categories or specific instances not encountered during training, placing higher demands on the model’s generalization and grounding capabilities. In object goal navigation, early end-to-end methods attempted to integrate textual knowledge into navigation tasks by leveraging compact multimodal feature spaces such as CLIP [53], yet their performance remains limited [31, 18, 47]. Modular approaches [26, 100, 1], on the other hand, typically rely on sensors for localization and mapping, high-level planning, and low-level control. Their dependence on hardware precision and pre-computation constrains practical flexibility. Meanwhile, instance goal navigation presents greater challenges due to its requirement for distinguishing unique target instances. Representative works such as Goat [9] achieve lifelong navigation by constructing an instance-aware semantic memory and integrating classical path planning. Mod-IIN [34], is specifically designed for this task, utilizing feature matching and map projection for goal re-localization. UniGoal [84] employs a unified graph representation and graph matching to handle diverse goal types. While these methods demonstrate strong performance on specific benchmarks, most still rely heavily on explicit environmental representations–such as maps or scene graphs–or extensive task-specific training. Our approach introduces an imagination-based, mapless navigation framework. This framework circumvents the need for extensive training by transforming the complex process of navigation planning into a selection problem based solely on RGB inputs.
Figure 2: The overall pipeline of our mapless, open-vocabulary navigation framework ImagineNav++. At each iteration, the agent captures a panoramic view of its surroundings. The Imagination Module then leverages the trained Where2Imagine module coupled with a novel view synthesis model to generate novel scene views. Guided by structured prompts, the VLM engages in target-oriented inference by integrating historical selective foveation memory with imagined future waypoint observations. Subsequently, the system executes the PointNav policy to determine the next navigational action. The above imagination, reasoning and planning procedure iterates until the target is reached.
@@ II-C Imagination in Embodied Navigation
Recent methods [90, 55, 102] have adopted supervised learning to infer potential target locations, addressing the ‘Where to look?’ challenge in navigation. These approaches typically predict either the absolute coordinates [90] of the target, the shortest distance to target [55], or the nearest boundary [102] based on local maps. Another line of research [54, 21, 41, 95] has focused on enhancing the prediction of unobserved regions through diverse methodologies. For instance, [54] introduced occupancy anticipation, where the agent infers an occupancy map based on RGB-D inputs. The L2M framework was introduced [21], consisting of a two-stage segmentation model that generates a semantic map beyond the agent’s field of view and selects long-term goals based on the uncertainty of predictions. SSCNav [41] leverages semantic scene completion and confidence maps to infer the environment and guide navigation decisions. A self-supervised generative map (SGM) is proposed [95], which employs self-supervised approach to continually generate unobserved regions in the local map and predict the target’s location. These methods primarily predict unobserved regions in top-down maps derived from egocentric RGB-D projections. By contrast, our ImagineNav++ performs future-scene imagination directly in RGB space. We employ a compact model aligned with human navigation behavior to generate novel viewpoint locations, which are then translated into corresponding visual observations through a diffusion-based view synthesis process.
@@ II-D Memory Mechanism in Embodied Navigation
Memory plays a crucial role in visual navigation by improving an agent’s perceptual consistency and long-term reasoning capabilities in unfamiliar environments [17, 71]. Initially, LLMs were predominantly applied in embodied navigation systems to support high-level decision-making. To enable LLMs to process perceptual inputs, earlier approaches [99, 93] generate descriptive captions of individual observations and aggregate these textual summaries over time. Subsequent methods typically construct a globally consistent memory map from visual observations, infer potential navigational waypoints through structured spatial reasoning, and encode these into natural language representations suitable for LLMs [63, 76, 36, 100], such as the semantic score map in OpenFMNav [36] and the reduced Voronoi graph in VoroNav [76]. The inherent loss of geometric information in language representations has motivated the development of hybrid memory approaches that maintain structured environmental state values (e.g., target distance and exploration reward). For instance, VLMap [26] quantifies navigation value via pixel-level feature-text similarity, while VLFM [85] projects image-text similarity into dynamic value maps using BLIP-2 [38]. While effectively mitigating memory forgetting, these methods require extensive precomputation and environment-specific training, limiting their scalability to unseen environments. With recent advances in multimodal learning, VLMs have been increasingly integrated into visual navigation systems to improve scene interpretation and decision-making [22, 65, 19, 92, 91, 69]. VLMNav [22] constructs dynamic traversability masks in a VQA framework. Concurrently, mapless navigation paradigms have emerged, bypassing explicit mapping to leverage multimodal observations. Representative methods include NaVid [92], which encodes monocular RGB streams into spatio-temporal representations using special tokens to differentiate historical and current observations; its extension Uni-NaVid [91] incorporating hierarchical token compression for efficient long-context processing; and the GRU-based history aggregation approach of [69], which models temporal observations to predict collision-aware action distributions. In line with these mapless navigation paradigms, our method also utilizes historical visual observations as direct input to VLMs. However, a key distinction of our approach lies in its use of visually salient keyframes, which compress perceptual history, enhance temporal coherence, and emphasize task-critical visual information, leading to more robust and efficient navigation performance.